The pizza palace offers three sizes of cola. The smallest size sells for $1.99, the medium for $2.49, and the large for $2.89. Thirty percent of the drinks sold are small, 50% are medium, and 20% are large. Create a probability distribution for the random variable price and answer the following questions.

a. Is this a discrete probability distribution? Indicate why or why not?

b. Compute the mean amount charged for a cola.

c. What is the variance in the amount charged for a cola? the standard deviation?
Answer #1

(a)

Let X is a random variable shows the prize of pizza. Following table shows the pdf of X:

X P(X=x)
1.99 0.3
2.49 0.5
2.89 0.2
Total 1

It is discrete random variable.

(b-c)

Following table shows the calculations for mean and SD:

X P(X=x) x*P(X=x) x^2*P(X=x)
1.99 0.3 0.597 1.18803
2.49 0.5 1.245 3.10005
2.89 0.2 0.578 1.67042
Total 1 2.42 5.9585

The mean amount charged for a cola is

E(X) = ΣχP(X = 1) = 2.42

The variance in the amount charged for a cola is
Var (X) = ΣπP(x = 1) - ΣπP(X = 1)] = 5.9585 – 2.42 0.1021

The standard deviation is

σ= Σ P(X = 1) - ΣπP(X = 1) = V5.9585 – 2.42 = 0.3195